OpenStudy (anonymous):

how would I distribute for: 1 radical 2 (3-radical 2)

OpenStudy (doc.brown):

By radical, do you mean \(\sqrt{x}\) ?

OpenStudy (anonymous):

yes

OpenStudy (doc.brown):

Do you mean\[1\sqrt{x}^2(3-\sqrt{x}^2)\]?